A local news team wants to estimate the number of people in a town of population 16,000, who have the Ring security system. They interview 250 homes. Of those polled, 116 had a Ring security system. How many homes in the city might one expect to have the security system? Round to the nearest whole number, if necessary.
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

116/250 * 16000 = 7424.

if the sample is a truly random sample of the homes in the population, then it is reasonable to expect that the proportion of homes that have a ring security system in the sample will be pretty close to the proportion of home that have a ring security system in the population.

116/250 = .464.

.464 * 16000 = 7424 plus or minus some number to account for the possibility of sampling error.










Answer by ikleyn(52873)

.

See the classic problem of the same type in the lesson
    - Using proportions to estimate the number of fish in a lake
in this site.

When you read this lesson, you will not miss this kind of problems with anything else.